Calculate the total and average kinetic energy of 32g of methane molecules at 27°C

Grade:11

1 Answers

Vikas TU
7 years ago
k.e = 3/2nRT
moles of methane = 32/16 = 2 mol.
put T = 300 K.
tis was total k.e
for avg. k.e. = 3/2 K T => 1.5 * K * 300
where K is the boltzman constant.